In Myanmar​, six ​laborers, each making the equivalent of $3.50 per​ day, can produce 40 units per day. In China, eleven ​laborers, each making the equivalent of $2.25 per​ day, can produce 45 units. In Billings, Montana, two ​laborers, each making $60.00 per​ day, can make 105 units.

Based on labor cost per unit​ only, the most economical location to produce the item is ______ (options are: billings/Montana, China, Myanmar), with a labor cost per unit of ​$______. ​(Enter your response rounded to two decimal​ places.)
Answer #1

Labor cost per unit in Myanmar = Number of laborers * Wage per day / Output per day

= 6*3.5/40

= $ 0.53 per unit

Labor cost per unit in China = 11*2.25/45

= $ 0.55 per unit

Labor cost per init in Billings/Montana = 2*60/105

= $ 1.14 per unit

Based on labor cost per unit​ only, the most economical location to produce the item is ___Myanmar___

(options are: billings/Montana, China, Myanmar), with a labor cost per unit of ​$__0.53____. ​(Enter your response rounded to two decimal​ places.)